# MSCPT: Few-shot Whole Slide Image Classification with Multi-scale and Context-focused Prompt Tuning

### v1

@@ IV-A3 Model Zoo
Thirteen influential approaches were employed for comparison, including traditional MIL-based methods: Mean pooling, Max pooling, ABMIL [32], CLAM [31], TransMIL [1], DSMIL [33] and RRT-MIL [40]; prompt tuning methods for natural images: CoOp [25], CoCoOp [24] and Metaprompt [26]; prompt tuning methods for WSIs: TOP [8] and ViLa-MIL[36]. Adapting to WSI-level tasks, we integrated an attention-based instance aggregation module [32] into the prompt tuning methods designed for natural images, such as CoOp, CoCoOp, and Metaprompt.
IV-A4 Implementation Details
@@
Following CLAM [31], the original WSIs were initially processed using the Otsu thresholding algorithm to remove the background parts. Subsequently, the WSIs were segmented into multiple non-overlapping patches of $256\times 256$ pixels at 5×5$\times$5 × and 20×20$\times$20 × magnification levels. We applied to perform our MSCPT on CLIP [14] and PLIP [22], both of which use ViT-B/16 [38] as their visual tower. Apart from MSCPT, Metaprompt, and DSMIL, which utilized inputs of both 5×5$\times$5 × and 20×20$\times$20 × magnification patches, the remaining methods solely relied on 20×20$\times$20 × magnification patches as inputs.
@@
For all methods, the Adam optimizer was employed with a learning rate of 1e-4, a weight decay of 1e-5 and batch size was set to 1. All methods were trained for a fixed number of epochs (100 for CLIP and 50 for PLIP) with early stop. We chose GPT-4 [41] to generate pathological visual descriptions, providing 10 low-level visual descriptions and 30 high-level visual descriptions for each category of WSIs (i.e., $C^{l}=10$ and $C^{h}=30$). For MSCPT and Metaprompt, we utilized the zero-shot capability of the VLMs to select 30 patches for each category at 5×5$\times$5 × magnification. The lengths of the global prompts $p_{glob}$ in both image and text encoder were uniformly set to 2. In this paper, unless explicitly stated otherwise, all experiments are conducted with 16 training samples per category. All the work was conducted using the PyTorch library on a workstation with eight NVIDIA A800 GPUs. All codes and details are released at https://github.com/Hanminghao/MSCPT.
IV-B Comparisons with State-of-the-Art
@@
The experimental results under the 16-shot setting are displayed in Table I. We observed some intriguing insights, such as complex and parameter-heavy methods like TransMIL and RRT-MIL underperformed despite their strong performance with full data training. Conversely, less parameterized methods such as ABMIL and CLAM exhibited slightly better performance. This is because traditional MIL-based methods require a lot of WSIs for training and the more parameters they have, the more training data is needed. Furthermore, after adapting the prompt tuning methods designed for natural images (i.e., CoOp, CoCoOp, and Metaprompt) to tasks at the WSI level, these methods outperform traditional MIL-based methods when based on CLIP and achieve comparable performance when using PLIP. Relatively few parameters contribute to this result. Additionally, we found that Metaprompt outperforms CoOp across most metrics, thanks to its integration of visual prompt tuning and multi-scale information. This result motivates us to pursue visual prompt tuning and develop more effective multi-scale information integration modules. Despite prompt tuning methods designed for Few-shot Weakly Supervised WSI Classification tasks having a relatively higher number of parameters, they exhibit the best performance. This is because VLMs’ prior knowledge is effectively exploited under the guidance of visual descriptive text prompts, alleviating the demand for extensive training data.
@@
Compared to other methods, our proposed MSCPT exhibits significant improvements in all evaluation metrics across the three datasets and two VLMs. Compared to the top-performing traditional MIL-based methods, MSCPT shows improvements of 0.3-13.0% in AUC, 2.0-12.3% in F1, and 1.5-11.6% in ACC across three datasets and two VLMs. Overall, MSCPT shows greater performance improvements when based on CLIP compared to PLIP. This is attributed to the specialized pre-training of PLIP on pathological images, enhancing its encoding capabilities for patches, thereby reducing the reliance on textual descriptions. Compared to the top-performing prompt tuning method suitable for natural images, MSCPT improved the AUC, F1, and ACC by 1.0-8.2%, 2.8-6.7%, and 1.6-6.9%.
TABLE II: Core components ablation experiment on the TCGA-NSCLC dataset based on PLIP.
@@
MHPT ISGPT NPCGP TCGA-NSCLC (PLIP-based)
@@
- - - 78.31±5.66 72.03±4.60 71.86±4.61
- - 80.57±4.17 73.62±2.41 73.77±2.48
- 82.75±5.73 75.48±4.70 75.53±4.72
- 81.92±5.03 74.96±4.68 75.10±4.64
84.29±3.97 76.39±5.69 76.54±5.49
TABLE III: Ablation experiment of different graph construction and training methods on the TCGA-RCC dataset based on CLIP.
Methods Trainable Param TCGA-RCC (CLIP-based)

### v2

@@ 4.1.3 Model Zoo
Thirteen influential approaches were employed for comparison, including traditional MIL-based methods: Mean pooling, Max pooling, ABMIL [29], CLAM [28], TransMIL [1], DSMIL [30] and RRT-MIL [40]; prompt tuning methods for natural images: CoOp [22], CoCoOp [21] and Metaprompt [23]; prompt tuning methods for WSIs: TOP [6] and ViLa-MIL [33]. Adapting to WSI-level tasks, we integrated an attention-based instance aggregation module [29] into the prompt tuning methods designed for natural images. Additionally, the prompts for CoOp, CoCoOp, and MetaPrompt were set according to the format provided in their official code: class name or “a photo of class name.
4.1.4 Implementation Details
@@
Following CLAM [28], the original WSIs were initially processed using the Otsu thresholding algorithm to remove the background parts. Subsequently, the WSIs were segmented into multiple non-overlapping patches of $256\times 256$ pixels at 5×5$\times$5 × and 20×20$\times$20 × magnification levels. We evaluate MSCPT on CLIP [12], PLIP [19], and CONCH [20], with the first two using ViT-B/32 [38] and CONCH using ViT-B/16 [38] as their visual towers. Apart from MSCPT, Metaprompt, and DSMIL, which used both 5×5$\times$5 × and 20×20$\times$20 × magnification patches, the remaining methods solely relied on 20×20$\times$20 × magnification patches as inputs. To ensure fair comparisons across different shot settings, 80% of each dataset was used for testing, while $K$ samples per class ($K=16,8,4,2$) were drawn from the remaining 20% for training.
@@
All methods used the Adam optimizer (learning rate: $41e-$ 4, weight decay: $51e-$ 5) with a batch size of 1. Training was fixed at 100 epochs for CLIP and 50 for PLIP and CONCH, with early stopping applied. We chose GPT-4 [43] to generate pathological visual descriptions, providing 10 low-level and 30 high-level visual descriptions for each category of WSIs (i.e., $C^{l}=10$ and $C^{h}=30$). For MSCPT and Metaprompt, we selected 30 patches for each category at 5×5$\times$5 × magnification for visual prompt tuning. The lengths of the global prompts $p_{glob}$ in both image and text encoders were uniformly set to 2. All experiments were conducted on a workstation with eight NVIDIA A800 GPUs.
Table 2: Results on five datasets, three downstream tasks under the 2-shot, 4-shot and 8-shot setting using CONCH. The highest performance is in bold, the second-best underlined, with mean and standard deviation reported across five random seeds.
{tblr}
colsep=1.5pt, rowsep = 0.05cm, column3 = c, column4 = c, column5 = c, column6 = c, column7 = c, column8 = c, column9 = c, column10 = c, column11 = c, column12 = c, column13 = c, column14 = c, column15 = c, column16 = c, column17 = c, column18 = c, column19 = c, column20 = c, column21 = c, column22 = c, column23 = c, row16 = bg=lightgray, row30 = bg=lightgray, row44 = bg=lightgray, cell11 = c=2,r=2c, cell13 = c=3, cell17 = c=3, cell111 = c=3, cell115 = c=3, cell119 = c=3, cell31 = r=14c, bg=lightgray, cell171 = r=14c, bg=lightgray, cell311 = r=14c, bg=lightgray, hline1,17,31,45 = -0.08em, hline2 = 3-5,7-9,11-13,15-17,19-21, hline3 = -, Methods & TCGA-NSCLC
(Subtyping, 2 classes) ~ TCGA-RCC
(Subtyping, 3 classes) ~ UBC-OCEAN
(Subtyping, 5 classes) ~ TCGA-BRCA
(Recurrence Prediction, 2 classes) ~ PANDA
(Grading, 3 classes)
AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC
2-shot 2 Training Samples per Class Max-pooling 80.85±5.19 64.68±6.49 67.31±4.22 86.62±5.10 69.08±8.05 73.21±7.63 84.82±2.76 58.70±5.87 62.65±5.75 67.92±6.34 61.66±7.31 67.60±5.07 70.31±3.67 51.20±4.20 56.44±3.67
Mean-pooling 78.29±5.24 70.35±4.44 70.79±4.32 93.08±2.79 77.45±6.52 80.63±5.33 85.69±2.97 57.82±5.36 59.66±8.76 66.34±8.16 58.96±8.70 62.57±7.07 71.59±5.06 51.57±4.66 56.66±5.60
ABMIL [29] 81.08±7.36 71.12±6.84 71.68±6.55 89.84±5.32 74.71±5.71 79.63±5.10 87.06±3.94 61.49±7.69 64.61±7.85 70.38±3.82 62.27±4.48 68.92±3.89 71.48±6.09 52.36±4.95 57.99±5.11
CLAM-SB [28] 83.73±5.36 72.67±3.61 73.25±3.26 93.05±3.01 79.99±6.52 83.99±5.75 88.11±4.40 62.13±4.53 62.84±4.14 67.45±8.24 60.25±7.22 64.73±8.64 71.16±6.91 50.69±9.72 54.50±12.12
CLAM-MB [28] 84.35±7.86 73.44±6.68 74.16±6.19 93.14±2.33 79.91±7.63 82.35±6.36 87.21±4.01 59.93±6.25 61.18±8.45 68.28±6.75 62.05±5.60 66.32±6.60 72.66±4.86 52.08±4.92 56.73±6.12
TransMIL [1] 78.97±5.95 71.34±5.45 71.63±5.36 93.70±2.84 76.99±1.52 79.74±5.88 86.64±3.58 59.42±5.64 61.23±7.68 64.51±4.71 60.88±6.05 67.30±5.04 70.53±6.26 49.80±5.97 54.90±6.86
DSMIL [30] 73.83±9.24 69.19±8.07 69.52±7.77 80.94±5.33 62.41±6.18 69.43±7.06 80.61±5.31 52.24±8.83 55.00±8.16 63.61±7.31 61.11±6.67 65.22±6.75 62.30±7.21 45.78±8.04 49.11±10.33
RRT-MIL [40] 80.74±7.70 74.37±6.25 74.40±6.22 92.03±2.55 77.33±6.77 79.60±6.26 85.75±3.78 57.92±3.49 58.04±5.82 68.57±5.50 62.48±6.47 69.08±2.24 70.09±8.40 48.42±6.94 54.93±6.57
CoOp [22] 76.93±6.08 69.93±5.21 70.12±5.14 91.62±3.38 75.58±6.00 78.14±5.29 82.78±5.85 53.18±8.62 55.00±10.26 67.71±5.68 59.62±4.55 64.25±4.63 72.08±6.49 49.33±6.59 56.03±5.80
CoCoOp [21] 79.63±5.73 71.87±5.44 72.19±5.25 93.06±2.78 77.79±6.88 80.77±6.02 85.62±1.97 59.68±3.12 63.28±4.47 66.17±7.28 62.89±6.71 66.33±7.46 72.30±4.68 49.31±2.67 53.66±6.24
Metaprompt [23] 87.11±4.37 78.30±5.21 78.51±4.99 94.26±2.58 82.45±4.68 84.58±4.12 85.86±2.89 58.80±5.86 60.64±7.29 67.61±4.87 61.78±4.86 67.35±5.57 75.09±4.60 53.43±3.45 58.44±3.99
TOP [6] 81.46±5.32 75.62±4.88 76.01±4.92 92.14±2.77 76.50±5.38 77.14±5.20 84.24±4.00 56.38±3.82 57.41±3.45 65.91±7.05 60.24±4.06 67.02±4.46 70.33±7.14 51.87±6.46 56.58±5.11
ViLa-MIL [33] 84.31±6.52 76.23±4.74 76.39±4.78 92.20±3.90 71.47±6.12 73.55±5.94 85.95±3.74 58.55±3.12 59.71±5.02 62.35±9.85 57.37±8.52 66.26±7.03 71.92±6.81 52.52±5.51 58.11±4.24
MSCPT(ours) 93.78±1.21 84.38±3.68 84.47±3.57 95.12±1.91 83.09±5.16 86.47±3.91 89.68±3.83 64.16±5.70 68.33±5.55 72.06±4.28 64.58±3.40 69.46±3.40 76.19±6.18 56.82±4.13 61.81±4.79
4-shot 4 Training Samples per Class Max-pooling 84.75±4.38 76.52±5.40 76.66±5.38 92.80±2.37 77.72±5.20 80.89±4.58 90.18±1.52 68.34±6.33 69.85±4.20 72.55±6.54 64.69±4.66 68.95±2.82 75.09±2.06 56.71±2.99 61.33±3.18
Mean-pooling 79.66±6.71 73.00±4.89 73.05±4.91 93.57±2.07 78.94±4.86 81.81±4.02 88.32±0.96 63.18±2.70 66.91±3.56 71.78±6.43 63.29±4.39 67.97±3.60 76.03±4.55 57.11±3.99 62.80±3.79
ABMIL [29] 84.54±7.61 75.71±7.44 75.89±7.30 93.01±4.97 81.29±4.20 84.56±3.22 91.72±2.13 69.60±7.34 72.79±7.44 74.59±6.82 64.16±6.97 68.87±6.51 79.17±4.97 59.48±3.45 64.98±4.87
CLAM-SB [28] 87.01±7.93 78.19±8.65 78.63±8.13 96.12±1.39 84.49±3.30 86.38±3.31 90.55±1.29 67.30±5.63 69.80±5.35 72.73±9.88 64.42±5.97 69.49±5.55 79.31±4.32 61.87±3.27 66.97±3.87
CLAM-MB [28] 86.19±10.61 78.71±10.68 78.92±10.44 96.24±2.11 84.69±4.83 86.93±4.27 92.04±1.41 70.72±4.94 72.55±4.95 73.20±7.79 65.08±5.97 71.15±3.26 79.63±4.13 60.80±3.68 65.32±4.55
TransMIL [1] 81.71±10.75 74.87±9.06 74.95±8.99 95.30±1.69 81.59±3.75 83.93±3.04 90.12±1.71 66.11±6.50 68.58±6.63 71.88±6.41 65.17±4.17 71.95±2.79 77.10±4.32 58.06±4.49 63.31±4.82
DSMIL [30] 76.42±12.49 72.49±9.98 72.62±10.00 92.00±5.12 77.37±9.20 82.52±6.56 85.84±7.61 60.82±13.07 62.99±14.10 67.63±6.40 63.63±6.65 68.44±5.05 74.37±5.24 55.33±3.78 61.85±4.99
RRT-MIL [40] 83.05±10.92 77.08±9.70 77.09±9.70 95.03±1.76 81.67±6.59 83.87±6.41 89.88±1.23 64.75±6.01 66.42±6.24 74.53±4.77 65.63±3.80 71.65±3.48 77.86±4.05 59.19±3.66 64.31±5.19
CoOp [22] 80.83±9.18 73.87±8.19 74.01±8.06 93.96±2.41 81.07±3.80 83.21±3.95 89.63±1.31 65.79±5.44 69.12±5.27 71.15±8.50 64.66±5.37 69.64±3.60 79.17±2.91 56.21±6.55 62.28±6.25
CoCoOp [21] 80.46±9.40 73.85±8.11 73.92±8.09 93.96±2.95 81.07±3.80 83.21±3.95 88.44±4.23 61.04±9.45 64.90±11.03 71.11±7.96 65.02±5.90 68.04±6.43 78.70±4.10 57.06±3.31 63.13±4.85
Metaprompt [23] 87.59±5.98 79.69±5.82 79.78±5.75 95.98±1.55 84.20±3.04 86.48±3.06 91.42±2.56 67.95±5.18 72.99±2.97 72.72±5.47
66.50±3.74 72.31±3.45 81.48±2.27 59.96±7.24 65.97±6.02
TOP [6] 84.49±7.66 77.50±7.51 77.53±7.32 95.01±1.44 82.46±3.87 84.70±3.61 90.45±1.11 67.89±4.05 69.04±3.72 72.65±6.10 65.66±4.97 68.18±5.78 78.43±4.72 58.44±2.19 64.39±2.76
ViLa-MIL [33] 85.67±8.20 78.15±8.19 78.20±8.13 94.65±1.72 79.24±4.67 82.09±4.74 89.72±0.90 63.88±3.86 66.23±3.24 70.21±8.80 62.72±5.51 68.44±6.17 76.01±5.10 57.83±3.56 62.85±3.60
MSCPT(ours) 93.60±2.76 86.13±3.58 86.15±3.54 97.03±2.05 87.26±5.03 89.48±4.02 93.19±1.97 73.53±5.26 76.81±5.10 76.20±4.17 68.25±4.82 73.60±3.00 82.33±3.51 64.13±1.51 68.41±2.46
8-shot 8 Training Samples per Class Max-pooling 91.54±1.81 83.41±3.23 83.46±3.16 97.39±0.54 88.47±1.53 90.17±1.10 91.72±1.15 71.43±1.55 73.84±2.08 74.62±1.91 66.56±1.69 69.70±2.10 77.49±4.34 59.38±4.25 64.19±4.15
Mean-pooling 85.37±2.94 77.42±3.52 77.52±3.41 96.61±0.38 86.26±1.27 88.34±1.18 91.83±1.15 70.48±2.87 72.79±3.43 72.47±2.84 64.63±2.16 68.65±1.90 78.07±4.34 60.21±5.46 64.49±3.92
ABMIL [29] 88.60±8.50 80.73±8.51 80.82±8.37 97.44±1.96 89.08±5.62 90.46±5.00 94.43±0.96 76.33±2.16 78.63±1.93 75.37±4.03 66.16±4.17 72.49±6.70 83.37±4.34 65.11±5.66 68.79±5.49
CLAM-SB [28] 92.62±5.11 85.92±5.76 85.94±5.73 97.56±0.28 90.87±0.88 91.12±0.59 94.44±0.74 77.12±1.51 78.22±1.96
77.02±4.66 67.96±4.58 73.02±3.11 83.93±2.82 65.67±4.42 69.67±3.95
CLAM-MB [28] 91.99±5.17 84.10±6.99 84.18±6.85 97.39±0.43 89.64±1.60 90.55±1.11 94.36±0.85 76.19±1.65 77.65±2.26 76.94±4.36 67.94±4.93
73.56±3.08 82.63±3.73 63.64±5.41 67.49±4.92
TransMIL [1] 88.71±2.96 80.81±3.94 80.91±3.82 97.97±0.58 89.78±1.09 91.32±0.93 93.53±0.62 72.97±2.27 75.05±3.31 75.39±2.61 67.52±2.16 72.17±2.39 80.68±4.10 62.45±5.10 66.92±4.77
DSMIL [30] 79.79±10.05 75.01±9.41 75.05±9.39 95.18±5.52 84.89±9.80 87.97±7.62 92.72±1.03 73.98±5.11 77.06±4.48 70.80±2.64 65.07±3.75 68.60±4.81 80.37±5.34 61.12±7.17 65.93±6.65
RRT-MIL [40] 90.82±3.45 84.27±3.83 84.33±3.82 97.42±1.18 88.06±3.37 89.86±2.74 93.43±1.21 72.92±4.47 75.00±4.81 76.17±4.62 66.77±4.29 73.01±3.49 80.56±4.78 61.50±5.66 66.40±5.26
CoOp [22] 87.36±4.42 79.43±5.30 79.57±5.16 97.26±0.64 88.19±1.39 89.63±1.66 92.98±0.72 72.62±2.89 75.54±3.39 76.03±5.25 66.94±3.60 70.17±3.11 81.22±2.19 61.56±5.81 67.55±5.29
CoCoOp [21] 88.61±2.71 80.26±4.38 80.38±4.23 97.36±0.55 88.39±0.91 89.94±1.16 92.57±1.94 72.10±4.05 75.39±3.94 76.68±3.67 67.73±3.16 73.22±3.21 81.74±2.16 63.58±2.97 67.82±2.39
Metaprompt [23] 91.54±3.39 83.74±5.07 83.85±4.91 98.05±0.32 90.47±1.19 91.81±0.83 92.99±2.03 74.94±2.51 78.14±2.73 76.88±2.02 68.02±2.39 72.95±1.07 82.67±1.73 65.21±4.46 69.17±4.42
TOP [6] 91.29±2.57 83.45±2.80 83.64±2.82 97.43±0.90 87.71±1.98 90.23±1.41 91.49±1.25 71.22±3.69 72.00±4.29 76.48±2.41 67.98±2.53 71.22±2.85 78.91±4.62 60.56±4.38 65.17±3.86
ViLa-MIL [33] 90.97±2.62 82.73±2.97 82.81±3.00 97.32±0.85 86.21±2.81 88.40±2.57 92.91±1.28 71.78±3.93 73.58±4.24 76.59±1.83 67.31±1.98 72.54±1.96 79.81±5.01 61.18±5.46 65.86±4.79
MSCPT(ours) 94.95±1.47 88.09±2.32 88.10±2.32 98.55±0.58 91.45±2.16 92.95±2.05 94.98±0.60 77.43±2.11 79.57±2.60 78.34±2.52 69.41±2.54 74.45±3.04 85.39±3.96 66.42±3.65 71.56±2.66
4.2 Comparisons with State-of-the-Art
@@
The experimental results under the 16-shot setting are displayed in Table 1. We observed some intriguing insights, such as complex and parameter-heavy methods like TransMIL and RRT-MIL underperformed despite their strong performance with full data training. Conversely, less parameterized methods such as ABMIL and CLAM exhibited slightly better performance. This is because traditional MIL-based methods require a large number of WSIs for training, and the more parameters they have, the more training data is needed. Furthermore, after adapting the prompt tuning methods designed for natural images (i.e., CoOp, CoCoOp, and Metaprompt) to tasks at the WSI-level, these methods outperform traditional MIL-based methods when based on CLIP and achieve comparable performance when using PLIP and CONCH. Relatively few parameters contribute to this result. Additionally, we found that Metaprompt outperforms CoOp across most metrics, thanks to its integration of visual prompt tuning and multi-scale information. This result motivates us to pursue visual prompt tuning and develop more effective multi-scale information integration modules. Despite prompt tuning methods designed for Few-shot Weakly Supervised WSI Classification (FSWC) tasks having a relatively higher number of parameters, they exhibit good performance. This leverages VLMs’ prior knowledge through visual and text prompt tuning, reducing the need for extensive training samples.
@@
Compared to other methods, MSCPT exhibits significant improvements in all evaluation metrics across the five datasets and three VLMs. Compared to the top traditional MIL-based methods, MSCPT shows the highest improvements in AUC, F1, and ACC by 13.0%, 12.3%, and 11.6%, respectively. Overall, MSCPT shows greater performance improvements when based on CLIP compared to PLIP and CONCH. This is due to PLIP and CONCH being pre-trained on pathological images, which enhances representation ability and reduces reliance on textual descriptions. Compared to the top prompt tuning methods for natural images (i.e., CoOp, CoCoOp, and Metaprompt), MSCPT achieves the highest improvements of 7.6%, 6.1%, and 6.5% in AUC, F1, and ACC, respectively.
Table 3: Ablation experiment of five core components.
@@
Prior Knowl. Multi Scale MHPT ISGPT NPCGP TCGA-NSCLC (PLIP-based, 16-shot) PANDA (CONCH-based, 4-shot)
AUC F1 ACC AUC F1 ACC
77.92±5.48 71.58±4.74 71.63±4.75 79.17±2.91 56.21±6.55 62.28±6.25
77.39±5.92 71.47±4.82 71.41±4.87 76.86±4.72 57.71±2.54 63.47±3.67
78.31±5.66 72.03±4.60 71.86±4.61 81.48±2.27 59.96±7.24 65.97±6.02
78.53±5.18 72.28±4.52 72.02±4.50 80.46±4.00 60.30±3.27 65.57±3.32
80.57±4.17 73.62±2.41 73.77±2.48 81.53±3.03 61.86±2.77 66.09±3.01
82.75±5.73 75.48±4.70 75.53±4.72 82.51±3.93 63.13±1.95 67.22±2.81
81.92±5.03 74.96±4.68 75.10±4.64 81.90±3.50 61.22±1.93 66.54±2.78
79.44±4.47 73.06±5.89 72.59±6.02 82.06±4.11 60.33±1.85 64.91±3.00
84.29±3.97 76.39±5.69 76.54±5.49 82.33±3.51 64.13±1.51 68.41±2.46
Table 4: Ablation experiment of Multi-scale Hierarchical Prompt Turning.
Global High Low UBC-OCEAN (CONCH-based, 4-shot)
@@
90.33±2.46 67.13±5.55 71.48±5.40
92.76±1.92 70.42±4.85 74.23±4.86
91.67±2.29 68.57±5.12 72.66±5.03
93.19±1.97 73.53±5.26 76.81±5.10
Prompt tuning methods explicitly designed for WSI exhibit superior performance on several datasets. This is attributed to their incorporation of priors into pre-trained VLMs and leveraging those priors to guide prompt tuning. Additionally, ViLa-MIL introduces multi-scale information compared to TOP, positioning it as the second-best overall performer. In comparison to ViLa-MIL, MSCPT shows improvements across all datasets, with AUC increasing by 0.5-5.3%, F1 by 2.4-7.4%, and ACC by 2.0-7.0%. This is because MSCPT performs prompt tuning on both the textual and visual modalities. Furthermore, MSCPT takes into account both the multi-scale and contextual information of WSIs. Unlike the late fusion approach in ViLa-MIL, MSCPT employs an intermediate fusion method, leveraging the transformer layer and trainable global prompts to integrate pathological visual descriptions from both high and low levels. Moreover, compared to the brief prior knowledge incorporated in ViLa-MIL, MSCPT introduces more detailed pathological visual descriptions. This enables MSCPT to achieve greater performance improvements over ViLa-MIL as the VLM becomes more advanced.
4.3 Results with Fewer Training Samples
To further validate MSCPT’s performance, we conducted experiments under 8-, 4-, and 2-shot settings. Notably, with limited training samples, sample selection critically influences model performance [6]. However, this influence diminishes as the performance of the VLM improves. Accordingly, this experiment was conducted based on CONCH, with the results presented in Table 2.
Overall, MSCPT consistently achieves the best performance across all datasets, with its relative advantage being most pronounced in low-data scenarios (e.g., 2-shot). Parameter-efficient methods such as CLAM and ABMIL demonstrate relatively stable performance, securing second-best results on certain datasets. In contrast, complex and parameter-heavy approaches like TransMIL and RRT-MIL perform poorly. Notably, MSCPT achieves significant improvements over the second-best model, with gains of up to 7.7% in AUC, 8.1% in F1, and 8.0% in ACC.
Table 5: Ablation experiment of different graph construction and training methods.
Methods Train. Param. TCGA-RCC (CLIP-based, 16-shot)